OpenStudy (anonymous):

y=10x^2-15x+5 & y=5x+5 find the point of intersection of the two systems. pls show me the working people.

OpenStudy (anonymous):

Equate the RHS of each equation\[10 x^2-15 x+5=5 x+5 \]and solve for x.\[10 (-2+x) x=0 \]x=0 and x=2 Use y = 5 x + 5 to find the corresponding y values. A plot is attached.

jhonyy9 (jhonyy9):

- so if you check those equations you see that for x=0 result that y=5 - so this (0,5) will be the point what you need to get